# itemizestditemize

### v1

@@ Abstract
We provide a distributed coordination paradigm that enables scalable and near-optimal joint motion planning among multiple robots. Our coordination paradigm is in contrast to the current paradigms that are either near-optimal but impractical for replanning times or real-time but offer no near-optimality guarantees. We are motivated by the future of collaborative mobile autonomy where distributed teams of robots will be coordinating via vehicle-to-vehicle (v2v) communication to execute information-heavy tasks such as mapping, surveillance, and target tracking. To enable rapid distributed coordination, we need to curtail the explosion of information-sharing across the network, thus, we need to limit how much the robots coordinate. However, limiting coordination can lead to suboptimal joint plans, causing non-coordinating robots to execute overlapping trajectories, instead of complementary. In this paper, we make theoretical and algorithmic contributions to characterize and balance this trade-off between decision speed and optimality. To this end, we introduce tools for distributed submodular optimization. Submodularity is a diminishing returns property typically arising in information-gathering tasks such as the aforementioned ones. On the theoretical side, we provide an analysis of how the network topology at the local level —each robot’s local coordination neighborhood— affects the near-optimality of coordination at the global level. On the algorithmic side, we provide a communication- and computation-efficient coordination algorithm that enables the agents to individually balance the trade-off. Our algorithm is up to two orders faster than competitive near-optimal algorithms. In simulations of surveillance tasks with up to 45 robots, the algorithm enables real-time planning at the order of $1$ Hz with superior coverage performance. To enable the simulations, we provide a high-fidelity simulator that extends AirSim by integrating a collaborative autonomy pipeline and simulating v2v communication delays.
@@
Multi-Robot Networks; Vehicle-to-Vehicle Communication; Submodular Optimization; Approximation Algorithms; Active Information Gathering
@@ I Introduction
In the future, distributed teams of robots will be coordinating via robot-to-robot communication to execute tasks such as collaborative mapping [1], surveillance [2], and target tracking [3] (Fig. 1). To this end, the robots will be adapting the information flow in the team to enable rapid and optimal decision-making (Fig. 2). Specifically, subject to their computation  and  communication bandwidth constraints,  the robots
@@
Figure 1: Scenario of Collaborative Road Detection and Coverage with 45 Aerial Robots. The drones are deployed in an unknown urban environment and tasked to collaboratively detect the roads and visually cover them. Particularly, the drones aim to collectively maximize the total covered road area in their top-view field of view (FOV). At each motion-planning step, each drone chooses a few others drones to coordinate with, subject to its onboard communication and computation bandwidth constraints. The full collaborative autonomy pipeline is depicted in Fig. 2. (a)–(c) depict the task progress across time: (a) The drones start from their initial deployment points; (b) Progress after 3 replanning steps; (c) The drones have achieved maximal road coverage.
will be choosing what information to receive and from who such that they can jointly plan actions in real-time, ensuring that their planned actions complement each other, instead of duplicating each other.
@@
Figure 2: Pipeline of Collaborative Mobile Autonomy. The robots sequentially perform in a distributed manner: (a) Network Self-Configuration: Given the observed environment and state of the robots, the robots decide with which other robots to communicate with, subject to their onboard bandwidth constraints; and (b) Action Coordination: The robots jointly plan actions —how to move and sense the environment— upon coordinating over the established communication network. (c) Action Execution: The robots execute their selected actions and perceive the environment. In this paper, our focus is on Action Coordination. Particularly, we present a communication- and computation-efficient distributed algorithm that is up to two orders faster than the state-of-the-art near-optimal algorithms. We also provide an analysis of how the network topology at the robot level —neighborhood of each robot— affects the near-optimality of coordination. The analysis offers insights into how the robots may configure the network to balance (tune) the trade-off between decision speed and near-optimality.
However, these capabilities of efficiency and effectiveness are currently challenging to achieve. The current literature on distributed coordination via vehicle-to-vehicle (v2v) communication imposes a trade-off between decision speed and optimality. On the one hand, algorithms that minimize the action overlap among robots may not be real-time since they can require an explosion of information sharing and processing across the robot network: processing and transmitting information cannot happen instantaneously —in the context of multi-robot tasks such as mapping, surveillance, and target tracking, real-time may indicate a replanning frequency at the order of $1$Hz. On the other hand, heuristic approaches can be real-time but cannot guarantee minimal action overlap since they achieve rapid decision-making by heuristically limiting the coordination among a few, possibly randomly chosen, robots. We need algorithms that can balance the trade-off by enabling the robots to choose what information to receive and from who.
Contributions. We show that it is possible to balance the trade-off of decision speed and near-optimality. Particularly, we provide a rigorous coordination approach that enables robots to self-configure their communication neighborhood (the set of robots to coordinate with) to tune the trade-off, that is, prioritize decision speed over near-optimality, as needed. To this end, we make theoretical and algorithmic contributions as follows, introducing tools for distributed submodular optimization. Submodularity is a diminishing returns property typically arising in information gathering tasks such as the aforementioned ones; we introduce it rigorously in Section III.
On the theoretical side, we provide an analysis of how the network topology at the local level —neighborhood of each robot— affects the near-optimality of coordination at the global level (Section V). The characterization quantifies the intuition that the more centralized the coordination (with larger neighborhoods), the slower the decision speed but the smaller the action overlap. It thus offers insights into how the robots can configure their network to balance the trade-off of decision speed and near-optimality. Importantly, we identify the existence of an inflection point that corresponds to the degree of centralization that optimizes the trade-off: with larger neighborhoods, we prove that the decision time increases faster than the action overlap decreases. Therefore, beyond a degree of centralization, the cost of sacrificing real-time performance will be higher than the gain in the actions.
On the algorithmic side, we present a communication- and computation-efficient distributed coordination algorithm that can be both real-time and near-optimal (Section IV). The algorithm is up to two orders faster than the competitive state-of-the-art near-optimal algorithms (Section VI). In simulated scenarios of map exploration with up to 45 robots, where we also simulate v2v communication delays (Section VII), the algorithm achieves replanning frequencies at the order of $1$Hz with superior performance. The algorithm’s decision time scales linearly with the number of robots, and sublinearly when parallelization is possible. For example, parallelization is possible in spatially distributed settings where robots are far enough from one another and have little action overlap.
The algorithm’s approximation guarantee captures the intuition that when a robot chooses not to coordinate with some other robots but ends up choosing an action that overlaps with these non-neighbors, then the achievable global optimality degrades. Specifically, we prove that the suboptimality degradation is proportional to the overlap. To enable our analysis, we quantify the suboptimality cost as a function of the action overlap between a robot and its (non-)neighbors. We introduce to this end a mutual-information-like quantity over sets that we term Centralization of Information.
The algorithm is built on a resource-aware distributed coordination framework introduced in Section III. The framework is resource-aware in that it requires each robot to coordinate actions with its neighbors only, receiving and processing information only about them instead of more robots in the network. As such, the framework is communication- and computation-efficient, curbing the explosion of information passing and processing in the network, thus, keeping low the delays due to limited computation and v2v communication speeds.
@@
Figure 3: Simulator pipeline. We provide a high-fidelity simulator that extends AirSim [4] to the multi-robot setting by integrating the autonomy pipeline of Fig. 2 and simulating v2v communication delays among other communication constraints. The illustrated pipeline is customized for the action coordination algorithm provided herein. The pipeline can be modified to other algorithms as desired. The code will be open sourced via a link in this paper.
Simulator. To demonstrate the efficiency and effectiveness of our coordination algorithm, we integrate it in a simulated state-of-the-art pipeline for active information-gathering with multiple robots (Fig. 3). To this end, we provide a high-fidelity simulator that extends AirSim [4] to the multi-robot setting, simulating v2v communication delays. Although the “vanilla” AirSim provides a simulation framework with a physics engine, flight controller, inertial sensors, and vision-based sensors, it provides no native support for multi-robot coordination that involves task orchestration and simulates no v2v communication delays. To enable such support, we develop appropriate frameworks and APIs. We will open-source the code upon the acceptance of the paper.
Evaluation. We tested the algorithm in the simulator over scenarios of collaborative mapping for road coverage with $15$ and $45$ robots (Fig. 1), respectively. Our algorithm yields practical advantages. In the case of $15$ robots, the algorithm achieves both superior road area coverage and real-time planning, at the order of $1$ Hz, being one to two times faster than the competitive state-of-the-art algorithms. In the case of $45$ robots, the robots are deployed into three spatially separated teams each of $15$ robots so parallelized decision-making is made possible. The algorithm demonstrates its ability to automatically scale with the decision time growing sublinearly. In more detail, it maintains superior road coverage and achieves similar planning time as in the case of $15$ robots, being at most 2 times slower. In contrast, the competitive state of the art is now 15 times slower.
@@
Organization of the Remaining Paper. In the following paragraphs, we compare with our preliminary work [5] and [6]. Then, in Section II, we present background on distributed coordination and related work. In Section III, we define the problem of Resource-Aware Distributed Coordination. Then, we present our algorithm in Section IV. The theoretical analysis of the algorithm’s approximation guarantees and decision time are presented in Sections V and VI, respectively. Section VII presents the simulator and the evaluation. Proofs not presented in the main text are presented in the appendix.
@@
Comparison with Preliminary Work [5] and [6]. This paper extends our preliminary work [5] to include new theory, the AirSim simulator implementation, the evaluations on the simulator, and proofs of all claims. Particularly, this paper provides novel performance guarantees, including an a posteriori bound with network-design implications, and an extension of the a priori bound in [5] to functions that are not submodular. Also, this paper bounds the decision time of the algorithm, characterizing for the first time its communication and computation complexity as a function of the size of the network and the times to perform computations and v2v communications. This paper introduces the AirSim-based simulator and the evaluations on the simulator. Instead, [5] employed MATLAB simulations only without simulating v2v communication delays. All proofs were omitted in [5], and here are presented for all original and new results.
@@
We also compare this paper with the pre-print that is under review for conference publication [6]. First of all, [6] introduces a different problem formulation from the one in eq. 1, requiring the communication-network topology to be co-optimized with the robots’ actions. The solution provided therein requires sensors/robots at fixed locations, instead of mobile sensors/robots, which are the focus herein. Finally, the results in [6] hold in probability and are based on regret optimization that requires quadratic time to converge, whereas the results herein are deterministic and are based on discrete optimization that requires linear time to converge.
II Background and Related Work
We first present background on distributed coordination, and explain that distributed coordination can be hard since it often takes the form of submodular set-function maximization, which is NP-hard. Then, we discuss the state of the art.
II-A Background on Distributed Coordination
Distributed coordination via v2v communication is challenging in multi-robot tasks such as mapping, surveillance, and target tracking, modeled as submodular set-function optimization problems [7, 2, 8, 9, 1, 3, 10, 11, 12, 13, 14], which are NP-hard [15, 16]. Submodularity is a diminishing returns property, capturing the intuition that when the same information is collected by two or more robots, its value to the task cannot be double-counted. Thereby, submodular multi-robot tasks are difficult to efficiently solve even in centralized settings. Rigorously, submodular multi-robot tasks take the optimization form:
$\max_{a_{i}\,\in\,\mathcal{V}_{i},\,\forall\,i\,\in\,{\cal N}}\ f(\,\{a_{i}\}_{i\,\in\,{\cal N}}\,),$ (1)
where ${\cal N}$ is the set of robots, $a_{i}$ is robot $i$’s action, $\{a_{i}\}_{i\,\in\,{\cal N}}$ is the ordered set of all robots’ actions, ${\cal V}_{i}$ is robot $i$’s set of available actions, and $f\colon 2^{\prod_{i\in{\cal N}}{\cal V}_{i}}\mapsto\mathbb{R}$ is the objective function that captures the (submodular) task utility.
For example, in road coverage with multiple drones (Fig. 1), ${\cal N}$ is the set of drones, ${\cal V}_{i}$ is the set of available trajectories the drone $i$ can choose from, and $f$ is the total road area covered by the drones’ collective field of view by the time their actions have been executed. This $f$ can be shown to be submodular. Intuitively, if two drones observe the same road area upon executing their actions, then either of the drones is redundant in the presence of the other. Therefore, to maximize the total covered road area, the drones need to minimize their action overlap via coordination.
Within the collaborative autonomy pipeline in Fig. 2, the optimization problem in eq. 1 models the action coordination stage, that is, the stage where robots jointly plan actions over a look-ahead horizon of optimization. Then, upon making the joint plans, each robot individually executes its action in a receding-horizon fashion till its action is updated via the subsequent coordination stage.
II-B Related Work
@@
We discuss the related work across four dimensions: (i) near-optimal but not necessarily real-time coordination algorithms; (ii) rapid but not necessarily near-optimal coordination algorithms; (iii) works on the trade-off between decision speed and optimality; and (iv) communication simulators.
Near-optimal coordination algorithms
@@
Although current approximation paradigms for the submodular maximization problem in eq. 1 [17, 16, 18] enable the robots to near-optimally minimize their action overlap [7, 2, 8, 9, 1, 3, 10, 11, 12, 13, 14], they may not be real-time. The reason is that their communication and computation complexities are superlinear in the number of robots, particularly, quadratic or cubic (we will elaborate on this later). Instead, we provide an algorithm with linear communication and computation complexity.
@@
Quadratic or higher communication and/or computation complexity can become prohibitive in large-scale networks due to real-world communication and computation delays. To illustrate the point, we give a toy example upon noting that communication delays are introduced by the limited v2v communication speeds. For example, state-of-the-art v2v communication speeds range from less than $1$Mbit/sec up to $125$MBytes/sec [19]. Computational delays are caused by the time required to perform function evaluation. This time depends on the processing required by the task at hand, e.g., image segmentation for the said task of road coverage (Fig 1). Assume now that the total delay per communication and computation is on average $1$msec. Then, for cubic decision-time complexity and $100$ robots ($|{\cal N}|=100$), the total time delay is at the order $100^{3}\cdot 1\text{msec}\simeq 17\text{min}$. This is three orders higher than the desired $\simeq 1Hz$ planning frequency.
@@
The quadratic or higher complexity of [17, 16, 18] over v2v networks is due to their coordination protocols: they instruct the robots to retain and relay information about all or most other robots in the network. Particularly, [17, 16] require iterative decision-making via consensus where at each iteration each robot needs to retain and transmit estimates of all robots’ actions [17, 16][18] requires sequential decision-making where all currently finalized actions need to be relayed to all robots in the network that have not finalized actions yet.
@@
In more detail, the multi-robot coordination algorithm in [20], inspired by [17, 16, 18], achieves the best possible approximation bound for the maximization problem in eq. 1, namely, $1-1/e$. However, it may require tenths of minutes to terminate in simulated tasks of $10$ robots even with no simulated communication delays [5]. The reason is that it requires a near-cubic number of iterations in the number of robots to converge (Table I). Similarly, for the Sequential Greedy (SG) algorithm [18], also known as Coordinate Descent, which is the gold standard in robotics and control for the maximization problem in eq. 1 [7, 2, 8, 3, 9, 1, 10, 11, 12, 13, 14], although it sacrifices some approximation performance to enable faster decision speed —achieving the bound $1/2$ instead of the bound $1-1/e$— still requires inter-robot messages that carry information about all the robots and, in the worst case, a quadratic number of communication rounds over directed networks [21, Proposition 2], resulting in the communication complexity being cubic in the number of robots [6, Appendix II].
Real-time coordination algorithms
@@
To achieve real-time planning, state-of-the-art coordination algorithms employ approaches that do not account for the submodular structure of the multi-robot task at hand. Thus, although fast, these approaches provide no guarantee that the agents’ joint actions will have minimal overlap. For example, in the context of collaborative exploration, the robots coordinate actions to move to the closest frontiers using shared map information in [22]. [23, 24, 25, 26] use an auction-based mechanism for task allocation among robots. In particular, the auction is also leveraged for high-probability communication maintenance in [25]. In [27], robots are assigned to different frontiers based on potential functions for distributed mapping and exploration. The robots set periodic meeting destinations in [28] to meet and coordinate actions during distributed exploration. [29] introduces a pair-wise coordination protocol where all pairs of neighboring robots sequentially coordinate trajectories to optimize their exploration task allocation. A similar pair-wise coordination protocol is proposed in [30] where each robot coordinates trajectories with the neighbor that it has longest not communicated with to minimize their overall trajectories.
Additionally, the said approaches are application-specific, e.g., exploration for the case of [28, 30], whereas, our theoretical and algorithmic results are general-purpose.
Characterization of the trade-off between decision speed and optimality
@@
To enable rapid distributed coordination at scale, we need to curtail the information explosion in multi-robot networks, thereby, need to limit what and how much information can travel across the network. However, the impact of imposing such information limitations is suboptimal actions.
@@
Current works have captured the suboptimality cost due to such information access for the case of the Sequential Greedy algorithm [31, 32, 12]. However, the provided characterizations are task-agnostic, considering the worst-case over all possible submodular functions and action sets.
In contrast, in this paper, the suboptimality cost is captured as a direct function of the action overlap over the task objective $f$ at hand and the robots’ chosen actions. Even if all robots plan in isolation, the guaranteed performance can still match the near-optimal $1/2$, if the robots’ chosen actions do not overlap, instead of being inversely proportional to the number of agents that would be the case for [31, 32, 12].
Also, to our knowledge, our characterizations are the first to quantify the suboptimality cost as a function of each agent’s local communication network of agents. The practical implication is that the characterizations can be used to enable the robots to select their communication neighborhoods to tune the trade-off of decision speed and near-optimality, subject to their communication bandwidth constraints.
Communication simulators for multi-robot applications
In information-heavy collaborative tasks such as mapping, surveillance, and target tracking, capturing how v2v communication delays can compromise the efficiency of multi-robot decision-making becomes important. Therefore, in this paper, given our focus on communication-efficient distributed decision-making, we indeed integrate in the AirSim pipeline v2v communication delays (Fig. 3).
For completeness, we next discuss papers that provide higher-fidelity simulations of wireless communication, including packet losses and occlusions. Particularly, [33] simulates the communication delay and demonstrates its influence in consensus-based applications. [34, 35, 36] simulate the communication channels and protocols and demonstrate how the communication delay between two robots will be impacted by the distance and line-of-sight conditions between them. [37] simulates realistic propagation models and scheduling functions for mobile radio-frequency (RF) communications and shows they can degrade task performance.
In our future work, we will integrate in the AirSim simulator such simulation capabilities to enable the testing of communication-aware and -efficient algorithms for distributed submodular coordination.
III Resource-Aware Distributed Coordination

### v2

@@ Abstract
We provide a communication- and computation-effi- cient method for distributed submodular optimization in robot mesh networks. Submodularity is a property of diminishing re- turns that arises in active information gathering such as mapping, surveillance, and target tracking. Our method, Resource-Aware distributed Greedy (RAG), introduces a new distributed optimization paradigm that enables scalable and near-optimal action coordination. To this end, RAG requires each robot to make decisions based only on information received from and about their neighbors. In contrast, the current paradigms allow the relay of information about all robots across the network. As a result, RAG’s decision-time scales linearly with the network size, while state-of-the-art near-optimal submodular optimization algorithms scale cubically. We also characterize how the designed mesh-network topology affects RAG’s approximation performance. Our analysis implies that sparser networks favor scalability without proportionally compromising approximation performance: while RAG’s decision time scales linearly with network size, the gain in approximation performance scales sublinearly. We demonstrate RAG’s performance in simulated scenarios of area detection with up to 45 robots, simulating realistic robot-to-robot (r2r) communication speeds such as the $0.25$ Mbps speed of the Digi XBee 3 Zigbee 3.0. In the simulations, RAG enables real-time planning, up to three orders of magnitude faster than competitive near-optimal algorithms, while also achieving superior mean coverage performance. To enable the simulations, we extend the high-fidelity and photo-realistic simulator AirSim by integrating a scalable collaborative autonomy pipeline to tens of robots and simulating r2r communication delays. Our code is available at https://github.com/UM-iRaL/Resource-Aware-Coordination-AirSim.
@@
Multi-Robot Mesh Networks; Robot-to-Robot Communication; Submodular Optimization; Approximation Algorithms; Active Information Gathering
@@ I Introduction
In the future, numerous distributed robots will be coordinating actions via robot-to-robot (r2r) communication to execute information-gathering tasks such as collaborative mapping [1], surveillance [2], and target tracking [3]. Such tasks require efficiency (scalability) and effectiveness (optimality), especially in crowded scenarios where the robots are many and operate close to each other; see, for example, the semantic-driven task of road detection with 45 aerial robots in Fig. 1.
But achieving scalability and optimality is hard. The first reason is that such tasks take the form of distributed submodular optimization, an NP-hard combinatorial optimization pro-
@@
Figure 1: Scenario of Collaborative Road Detection with 45 Aerial Robots. The drones are deployed in an unknown environment and tasked to collaboratively detect the roads and visually cover them. Particularly, the drones aim to collectively maximize the total covered road area in their top-view field of view (FOV). At each motion-planning step, each drone chooses a few other drones to coordinate with, subject to its onboard communication and computation bandwidth constraints. The full collaborative autonomy pipeline is depicted in Fig. 2. (a)–(c) depict the task progress across time: (a) The drones start from their initial deployment points; (b) Progress after 3 replanning steps; (c) The drones have achieved maximal road coverage.
@@
Figure 2: Pipeline of Collaborative Mobile Autonomy. The robots sequentially perform over a mesh network: (a) Network Self-Configuration Step: Given the observed environment and state of the robots, the robots decide with which other robots to communicate with, subject to their onboard bandwidth constraints; and (b) Action Coordination Step: The robots jointly plan actions —how to move and sense the environment— upon coordinating over the established communication network. (c) Action Execution Step: The robots execute their selected actions and perceive the environment. In this paper, our focus is on Action Coordination over tasks that take the form of submodular optimization. We present a communication- and computation-efficient distributed algorithm that scales linearly with the network size, in contrast to the cubic time complexity of the competitive near-optimal algorithms. Along with our algorithmic contribution, we provide a rigorous analysis of how each agent’s communication neighborhood affects the near-optimality of the optimization. The analysis implies that establishing sparser neighborhoods favors scalability without proportionally compromising approximation performance.
blem [4]. Therefore, such tasks require increased computations and communications to be solved optimally. Submodularity is a property of diminishing returns, and is encountered across robotics [5] as well as machine learning [6] and control [7]. It captures the intuition that when any two robots are nearby and collect same information, e.g., track the same targets, then one of the robots is redundant (it would be more effective if the robots were collecting different information). On top of the above reason, another fundamental reason is the communication and computation limitations of the robots versus the r2r messaging requirements by the tasks. For example, in active distributed simultaneous localization and mapping (SLAM), the inter-robot communication messages can grow to a few MBs within a few hundred meters of explored environment [8] ($1$ MB $=8$ Mb). But the wireless r2r communication speeds range from 0.25 Mbps, e.g., achieved by the Digi XBee 3 Zigbee 3.0 antenna module, to 100 Mbps, e.g., achieved by the Silvus Tech SL5200 antenna module. Therefore, transmitting even a single message can take seconds.
The current approaches are either near-optimal but not scalable, or real-time but offer no performance guarantees. The near-optimal methods [9, 10, 11] that are being used for active information gathering with multiple robots [12, 2, 13, 14, 1, 15, 16, 17, 18, 19, 20], scale cubically with the network size, resulting in impractical running times in mesh networks with tens of robots (up to tens of minutes per action coordination step in the presence of real-world communication delays). The state-of-the-art multi-robot motion planning methods are real-time [21, 22, 23, 24] but they are suitable in spatial exploration where robots are expected to operate far from each other. Then, the submodular structure can be ignored and opportunistic coordination is sufficient. We discuss the related work in extension in Section II.
@@
Figure 3: Simulator pipeline. We provide a high-fidelity simulator that extends AirSim [25] to the multi-robot setting by integrating the autonomy pipeline of Fig. 2 and simulating r2r communication delays among other communication constraints. The illustrated pipeline is customized for the action coordination algorithm provided herein. The pipeline can be modified to other algorithms as desired.
In this paper, instead, we focus on crowded scenarios where robots operate close to one another, and their objective is not merely exploration; it is discovery and/or monitoring of specific information, e.g., landmarks, mobile targets, areas and objects with specific semantics [1, 26, 5] (Fig. 1). Then, the submodular structure is present. Thus, to jointly maximize the submodular objective, the robots must intelligently coordinate actions to enable both scalability and near-optimality.
Contributions. We provide a communication- and computa- tion-efficient method for distributed submodular optimization in robot mesh networks. The method, Resource-Aware distributed Greedy (RAG), enables both scalable and near-optimal coordination over robot mesh networks despite real-world communication and computational delays. In contrast, the current approaches are either near-optimal but not scalable or are real-time but offer no performance guarantees. RAG applies to any distributed submodular optimization task. In this paper, we evaluate RAG in a semantics-driven area exploration task of road detection and coverage with tens of robots (Fig. 1), via high-fidelity simulations. In more detail, our technical and simulator contributions are as follows.
We characterize RAG’s time complexity and approximation performance as a function of the state of the robot mesh network such as the robots’ communication neighborhoods, task objective function, and expected delays due to the robots’ computation and communication capabilities:
  • Time complexity: RAG’s decision-time scales linearly with the network size. Instead, the competitive near-optimal submodular optimization algorithms scale cubically with the network size. To this end, RAG introduces a resource-minimal distributed optimization paradigm that requires each robot to make decisions based only on locally observed information and information received from and about their neighbors only. In contrast, the current submodular optimization paradigms [9, 10, 11], widely applied in robotics and control [12, 2, 13, 14, 1, 15, 16, 17, 18, 19, 20], that instead require the relay of information about all robots across the network.
  • Approximation performance: We present suboptimality bounds for RAG. The bounds are useful as follows: (i) They capture the approximation performance of RAG as a function of each agent’s local mesh-network topology, quantifying how denser mesh networks (more coordination) improve the approximation performance. (ii) The bounds inform how to design each agent’s communication neighborhood to balance the trade-off between decision speed and optimality. Particularly, we prove that the approximation performance gain grows sublinearly with the agents’ communication neighborhood size. Therefore, in combination with the fact that RAG’s decision time scales linearly with the network size, the bounds imply that sparser neighborhoods would favor decision speed without proportionally compromising approximation performance. The appropriate level of neighborhood sparsity that balances the trade-off for the task at hand can be found via experimentation in high-fidelity simulations, as we illustrate in our simulations.
To perform high-fidelity evaluations, we enhance the physics-based and photo-realistic simulator AirSim [25] to enable scalable simulation of robot mesh networks, including simulation of r2r communication delays (Fig. 3). To this end, we first implement multi-threading for sensor data processing in ROS wrappers, resulting in a tenfold improvement in data frequency across multiple sensing modalities, including RGB, depth, segmentation, IMU, GPS, and barometer. Second, we develop an r2r coordination module that models realistic communication delays and ranges between drones, with signal attenuation and dynamic neighborhood relationships managed through the ROS master. Finally, we implemented automatons for capturing segmentation images via rotating gimbals and calculating marginal information gain by stitching semantic data from multiple drones. The implementation uses ROS1, thus inherently incorporating its transport layer delays and message passing overhead, as shown in Figs. 710.
Evaluations. We evaluate RAG in the provided simulator across four scenarios of road detection and coverage (Fig. 1). The scenarios include two team sizes, 15 and 45 robots, and two communication rates, 0.25 Mbps, e.g., achieved by the Digi XBee 3 Zigbee 3.0 antenna module, and 100 Mbps, e.g., achieved by the Silvus Tech SL5200 antenna module.
RAG achieves real-time planning, up to three orders of magnitude faster than competitive near-optimal submodular optimization algorithms; and, when the robots maintain a neighborhood size $\geq 2$, RAG also achieves superior mean coverage. The experiments also demonstrate RAG scales linearly with the network size, as predicted by our theoretical analysis: from the 15-robot to the 45-robot case, RAG scales linearly, being at most 3x slower compared to the 15-robot case. In contrast, the compared near-optimal algorithms scale cubically, being up to 60x slower compared to the 15-robot case.
@@
Organization of the Remaining Paper. In Section II, we present background on distributed submodular coordination and related work. In Section III, we define the problem of Resource-Aware Distributed Submodular Optimization. Then, we present our algorithm in Section IV. The theoretical analysis of the algorithm’s approximation guarantees and decision time are presented in Sections V and VI, respectively. Section VII presents the simulator and the evaluations. All proofs are presented in the Appendix.
@@
Comparison with Preliminary Work [27] and [28]. This paper extends our preliminary work [27] to include new theory, the AirSim extension, the evaluations, and proofs of all claims. Particularly, this paper provides novel performance guarantees, including an a posteriori bound with network-design implications, and an extension of the a priori bound in [27] to functions that are not submodular. Also, this paper bounds the decision time of the algorithm, characterizing for the first time its communication and computation complexity as a function of the network size and the time to perform computations and r2r communications. This paper introduces the AirSim-based simulator and the evaluations on the simulator. Instead, [27] employed MATLAB simulations only without simulating r2r communication delays. All proofs were omitted in [27], and here are presented for all original and new results.
@@
We also compare this paper with our preliminary work [28]. First of all, [28] introduces a more complex problem formulation that requires the mesh-network topology to be co-optimized with the robots’ actions. The solution provided therein requires robots at fixed locations, instead of mobile sensors, which are the focus herein. Finally, the results in [28] hold in probability and are based on regret optimization that requires quadratic time to converge, whereas the results herein are deterministic and are based on discrete optimization that requires linear time to converge.
II Related Work
@@
We discuss related work on (i) near-optimal but not necessarily real-time distributed submodular optimization, (ii) rapid but not necessarily near-optimal distributed multi-robot motion planning, (iii) the trade-off of decision speed vs. optimality, and (iv) high-fidelity communication simulators.
Near-optimal distributed submodular optimization
@@
The current approximation paradigms for submodular maximization problem [9, 10, 11] have been widely used in active information gathering with multiple robots [12, 2, 13, 14, 1, 15, 16, 17, 18, 19, 20] (see also the survey [5]). However, they may not scale. The reason is that their communication and computation complexities are superlinear in the number of robots, particularly, quadratic or cubic (Table I). Instead, we provide an algorithm with linear communication and computation complexity.
@@
Quadratic or higher communication and/or computation complexity can become prohibitive in large-scale networks due to real-world communication and computation delays. To illustrate the point, we give a toy example upon noting that communication delays are introduced by the limited r2r communication speeds. For example, state-of-the-art r2r communication speeds range from less than $1$Mbps up to $100$Mbps [29]. Computational delays are caused by the time required to perform function evaluation. This time depends on the processing required by the task at hand, e.g., image segmentation for the said task of road coverage (Fig 1). Assume now that the total delay per communication and computation is on average $1$msec. Then, for cubic decision-time complexity and $100$ robots ($|{\cal N}|=100$), the total time delay is at the order $100^{3}\cdot 1\text{msec}\simeq 17\text{min}$.
@@
The quadratic or higher complexity of [9, 10, 11] over r2r networks is due to their coordination protocols: they instruct the robots to retain and relay information about all or most other robots in the network. Particularly, [9, 10] require iterative decision-making via consensus where at each iteration each robot needs to retain and transmit estimates of all robots’ actions [9, 10][11] requires sequential decision-making where all currently finalized actions need to be relayed to all robots in the network that have not finalized actions yet.
@@
In more detail, the multi-robot coordination algorithm in [30], inspired by [9, 10, 11], achieves the best possible approximation bound for submodular maximization, namely, $1-1/e$. However, it may require tens of minutes to terminate in simulated tasks of $10$ robots even with no simulated communication delays [27]. The reason is that it requires a near-cubic number of iterations in the number of robots to converge (Table I). Similarly, for the Sequential Greedy (SG) algorithm [11], also known as Coordinate Descent, which is the gold standard in robotics and control for submodular task optimization [12, 2, 13, 15, 14, 1, 16, 17, 18, 19, 20], although it sacrifices some approximation performance to enable faster decision speed —achieving the bound $1/2$ instead of the bound $1-1/e$— has communication complexity that is cubic in the number of robots [28, Appendix II]. The reason is that it requires inter-robot messages that carry information about all the robots and a quadratic number of communication rounds over directed networks [31, Proposition 2].
Real-time multi-robot motion planning
@@
Current motion planning methods also employ approaches that do not need to account for the submodular structure of the multi-robot task: the robots are expected to operate far from one another, collecting decoupled information. Then, the submodular objective simplifies to a distributed, modular objective: it can be expressed as the sum of decoupled functions, one for each robot, each computable by the robot based on its action only. Therefore, these methods often allocate robots to unexplored areas and instruct the robots to explore them independently. Opportunistic communication is sufficient, when the robots happen to be nearby. For example, in collaborative exploration, the robots coordinate actions to move to the closest frontiers using shared map information in [32]. [33, 34, 35, 21] use an auction-based mechanism for task allocation among robots. The auction is also leveraged for high-probability communication maintenance in [35]. In [22], robots are assigned to different frontiers based on potential functions for distributed mapping and exploration. The robots set periodic meeting destinations in [23] to meet and coordinate actions during distributed exploration. [36] introduces a pair-wise coordination protocol where all pairs of neighboring robots sequentially coordinate trajectories to optimize their exploration task allocation. A similar pair-wise protocol is proposed in [24] where each robot coordinates trajectories with the neighbor that it has longest not communicated with.
In contrast, this paper focuses on crowded scenarios where robots operate close to one another, and their objective is not merely exploration; it is discovery and/or monitoring of specific information, e.g., landmarks, mobile targets, and areas or objects with specific semantics [1, 26, 5]. Thus, to jointly maximize the submodular objective, the robots must coordinate actions [1]: the submodular objective cannot be decomposed as the sum of locally computable objectives, one for each robot, where the actions of each robot are unaffected by the actions of other robots. Additionally, while the discussed task allocation works are application-specific, e.g., exploration [23, 24], our algorithmic results are general-purpose and apply to any distributed submodular optimization setting in robotics, control, machine learning, and beyond.
Trade-off of decision speed vs. optimality
@@
To enable rapid distributed submodular optimization, we need to curtail the information explosion in robot mesh networks. Thereby, we need to limit what and how much information can travel across the network. However, the consequence of imposing such information limitations is suboptimal coordinated actions.
@@
Current works have captured the suboptimality cost due to such limited information access for the case of the Sequential Greedy algorithm [37, 38, 18]. The provided characterizations are task-agnostic, holding true for the worst case over all possible submodular functions and action sets. In contrast, we capture the suboptimality cost as a function of the task objective $f$ at hand, the robots’ coordinated actions, and the current mesh network topology. Thus, our characterizations, being task specific, (i) can be tighter, and (ii) can be used by the robots to design communication neighborhoods that tune the trade-off of decision speed and near-optimality, subject to their communication bandwidth constraints.
Communication simulators
Papers focus on high-fi- delity simulations of wireless communications, including the simulation of protocols and communication limitations such as communication delays and package dropouts. [39] simulates communication delays, and demonstrates its influence in consensus-based applications. [40, 41, 42] simulate communication channels and protocols, analyzing how distance and line-of-sight conditions impact communication delays. [43] simulates realistic propagation models and scheduling functions for mobile radio-frequency communications. We instead focus on enabling large-scale photo-realistic simulations with tens of robots, analyzing how communication delays impact active information gathering with robot mesh networks.
III Resource-Aware Distributed Submodular Optimization